The ratio of the complement of an angle to the supplement of the angle is 7:22. Find the measure of th angle, the complement, and the supplement.
Answer by rfer(16322)

90-7x=180-22x
15x=90
x=6
90-7x=48 compliment=42
180-22x=48 supplement=132